OpenStudy (anonymous):

What is the total charge of 1.0g of electrons?

OpenStudy (anonymous):

1.0 gram (g) of electrons would contain 10^27 electrons Electrons have an electric charge of −1.602×10−19 coulomb so total charge of 1 g electrons = -1.602 x 10^-19 x 10^27 = -1.602 x 10^8

OpenStudy (anonymous):

mass of 1e=9.10938291(40)×10^-31kg =9.109*10^-28g charge mass −1.602176565(35)*10^-19 9.109*10^-31kg X 9.109*10^-28g => X=9.109*10^-31 * 9.109*10^-28g ------------------------------- −1.602*10^-19 =>5.179393320849e-77 i have done my best sorry if wrong

OpenStudy (anonymous):

One electron mass is 9.109*10^(-28) g, so there are 1/(9.11*10^(-28)) = 1.098*10^27 electrons per gram. One electron charge is 1.602*10^(-19) C, so the charge of a gram of electrons is (1.098*10^27)(1.602*10^(-19)) C = 1.759*10^8 C.